OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5). Answer 52 22 -1.5 102

OpenStudy (anonymous):

f(5)=2(5)^2+2 =2(25)+2 =50+2 =52

OpenStudy (anonymous):

the ^ sign is usually used to denothe that something is being rasied to a power